The ages of a group of

156156

randomly selected adult females have a standard deviation of

16.516.5

years. Assume that the ages of female statistics students have less variation than ages of females in the general​ population, so let

sigmaσequals=16.516.5

years for the sample size calculation. How many female statistics student ages must be obtained in order to estimate the mean age of all female statistics​ students? Assume that we want

9898​%

confidence that the sample mean is within​ one-half year of the population mean. Does it seem reasonable to assume that the ages of female statistics students have less variation than ages of females in the general​ population?
Answer #1

Solution :

Given that,

Population standard deviation = \sigma = 16.5

Margin of error = E = 0.5

At 98% confidence level the z is ,

\alpha = 1 - 98% = 1 - 0.98 = 0.02

\alpha / 2 = 0.02 / 2 = 0.01

Z\alpha/2 = Z0.01 = 2.326

sample size = n = (Z\alpha/2* \sigma / E) 2

n = (2.326 * 16.5 / 0.5)2

n = 5891.79

n = 5892

Sample size = 5892